Required information [The following information applies to the questions displayed below.] On January 1, year 1, Dave received 1,150 shares of restricted stock from his employer, RRK Corporation. On that date, the stock price was $8 per share. On receiving the restricted stock, Dave made the 83(b) election. Dave’s restricted shares will vest at the end of year 2. He intends to hold the shares until the end of year 4 when he intends to sell them to help fund the purchase of a new home. Dave predicts the share price of RRK will be $38 per share when his shares vest and will be $42 per share when he sells them. Assume that Dave’s price predictions are correct and answer the following questions: (Leave no answers blank. Enter zero if applicable. Round your final answer to the nearest whole dollar value. Enter all amounts as positive values.) a. What are Dave’s taxes due if his ordinary marginal rate is 32 percent and his long-term capital gains rate is 15 percent?
Answer #1

Taxes due Formula Grant Date Vesting pate $11,776 = (32* 1,150J* 32% sale date $ 1,728 = (42* 1,150 - 32*1150) 15% Tax on gra